# Open Materials Generation with Stochastic Interpolants

### v1

@@ III.3 Joint generation with stochastic interpolants
For both CSP and DNG tasks, we seek to generate samples from a joint distribution over multiple coordinates. For DNG, this joint distribution $\rho_{1}$ encompasses all elements of a crystal unit cell. For CSP we similarly model the joint distribution, $\rho_{1}$, but with atom types fixed to compositions sampled from the target dataset. For both tasks, the total loss function is formulated as a weighted sum of the individual loss functions for each variable (see Appendix B.2), and their relative weights are optimized (see Appendix B.3). We illustrate both types of models and their structure generation process in Fig. 2.
Figure 2: Illustration of CSP and DNG tasks. (a) For CSP, the species $\bm{A}$ are fixed with known compositions from $t=0$. From this, we predict $\bm{X}$ and $\bm{L}$ from randomly sampled initial values. For DNG, we predict $(\bm{A},\bm{X},\bm{L})$ jointly. Our implementation of DFM initializes $\bm{A}$ as a sequence of masked particles that are unmasked through a series of discrete jumps to reveal a physically reasonable composition. (b) Two avenues for performing de novo generation of materials. The first uses two steps: a CFP model predicts compositions and then uses a CSP model to find accompanying stable structures. The second trains a DNG model over cell, species, and fractional coordinates jointly as shown in (a).
Additionally, for DNG, we consider a two-step process in which composition is learned separately from structure, as seen in Fig. 2b. In this approach, we first train a Chemical Formula Prediction (CFP) model (see Appendix B.1) to generate compositions optimized for SMACT stability [43], similarity in the distribution of $N$-arity of known structures, as well as uniqueness and novelty. The predicted compositions are then used as input for a pretrained CSP model, which generates the corresponding atomic configurations.
III.4 Data-dependent coupling
SIs have been used with data-dependent couplings [44], where a coupling function $\nu(x_{0},x_{1})$ enables biasing of $x_{0}$ based on the sampled $x_{1}$. In OMG, we incorporate an optional data-dependent coupling that enforces an ordering (i.e., a permutation on the order of atomic elements within a structure) that produces the minimum fractional-coordinate distance between each particle pair $(\bm{x}_{0}^{i},\bm{x}_{1}^{i})$ from structures $x_{0}\in\rho_{0}$ and $x_{1}\in\rho_{1}$ (see Appendix A.6). We find that the inclusion of this data-dependent coupling is optimal during hyperparameter tuning depending on the type of model: CSP models typically performed better without this coupling, but DNG models (see Tab. 4) can benefit in certain cases from minimizing traveled distance via permutation of elements.
III.5 Velocity annealing
Velocity annealing—rescaling the learned drift during generation to increase velocity over time as $b^{\theta}(t,x)\rightarrow(1+st)\,b^{\theta}(t,x)$ with $s$ as an hyperparameter during integration—has been empirically shown to improve performance in a number of studies that apply CFM to physical systems [45, 46, 32]. For instance, Miller et al. [32] demonstrated that applying velocity annealing significantly improves performance in CSP and DNG benchmarks for materials. Motivated by these findings, we include velocity annealing in OMG as a tunable hyperparameter, while emphasizing that this technique lacks a formal theoretical justification within the mathematical frameworks underlying flow models and stochastic interpolants.
@@ IV.1 Choice of interpolant
In training OMG, we optimize the choice of the interpolating function that is used during training. Albergo et al. [33] introduced four interpolants of the form defined in Eq. 1, each shaping the interpolation trajectory differently (see also Appendix A.2 for further details).
@@
Table 1: Results from crystal structure prediction. Match rate and RMSE of matched structures without (left) and with (right) filtering for structural and compositional validity are reported for all models. For OMG’s model, the choice of positional interpolant, latent variable component $\gamma$, and sampling scheme are noted.
@@
OMG 80.85/79.55 0.3864/0.3873 69.83/63.75 0.0741/0.0720 27.38/25.15 0.1970/0.1931 69.44/61.38 0.1303/0.1289
Trig; ODE; $\gamma=\sqrt{0.113t(t-1)}$ Linear; ODE; $\gamma=0$ Linear; ODE; $\gamma=0$ Linear; ODE; $\gamma=0$
The linear interpolant defines a constant velocity trajectory from $x_{0}$ to $x_{1}$. When combined with an ODE sampling scheme and $\gamma=0$, this reproduces a particular choice of CFM. If the sampling scheme or $\gamma$ are changed, however, the use of linear interpolants become distinct from CFM. The trigonometric interpolant prescribes trajectories with more curvature than the linear interpolant. The encoder-decoder interpolant first evolves samples from $\rho_{0}$ to follow an intermediate standard Gaussian distribution $\mathcal{N}(0,\bm{I})$ at time $t=0.5$, before mapping them to $\rho_{1}$ at $t=1$. This approach has been found to interpolate more smoothly between distributions, potentially mitigating the formation of spurious features in the probability path at intermediate times [33]. Lastly, we consider the variance preserving score-based diffusion (SBD) interpolant. When paired with an SDE sampling scheme, this interpolant is mathematically equivalent to an SBDM, but on the continuous time interval $[0,1]$. The SBD interpolant assumes that $\rho_{0}$ is a Gaussian, and unlike the previous three interpolants it involves no explicit latent variable; instead the $\alpha(t)x_{0}$ term takes on this role. We note that the trajectory of the encoder-decoder interpolant between times $t=0.5$ and $t=1$ resembles that of the SBD interpolant between times $t=0$ and $t=1$. For the example of using the encoder-decoder interpolant only for the coordinates, however, we emphasize that the Gaussian-distributed coordinates at $t=0.5$ are conditioned on other coordinates that are halfway-interpolated at this point. Conversely, for SBD interpolation, the Gaussian distributed coordinates at $t=0$ are only conditioned on other random variables since, at this point, all elements of $x_{0}$ are randomly distributed.
We prioritize flexibility in learning the atomic coordinates, $\bm{X}$, as our findings indicate that learning drifts and denoisers which generate the positional distribution with high fidelity are more difficult than for the other two coordinates of interest. As such, we evaluate the periodic versions of the four interpolants outlined above on learning the generation of atomic positions $\bm{X}$, but only use the linear interpolant with $\gamma=0$ for the lattice vectors $\bm{L}$.
IV.2 Equivariant representation of crystal structures
In the current OMG implementation, we employ CSPNet as an encoder that is trained from scratch. The CSPNet architecture encodes atomic types using learnable atomic embeddings and represents fractional coordinates through sinusoidal positional encodings (see Appendix B.1). These features are processed through six layers of message-passing, after which the encoder produces the drift $b^{\theta}(t,x)$ of both the lattice and the fractional coordinates, as well as potentially predicting the denoiser $z^{\theta}(t,x)$. For DNG, the network must also predict $\log p_{1|t}^{\theta}(\bm{a}_{1}|x_{t})$. The resulting outputs inherently preserve the permutation, rotational, and translational symmetries embedded in CSPNet.
The output of CSPNet is invariant with respect to translations of the fractional coordinates in the input. Thus, one should, in principle, use a representation of the fractional coordinates that does not contain any information about translations. While this is straightforward in Euclidean space by removing the mean of the coordinates of the given structure, this cannot be done with periodic boundary conditions where the mean is not uniquely defined. We follow Miller et al. [32] and instead remove the center-of-mass motion when computing the ground-truth $\partial_{t}x(t,x_{0},x_{1},z)$ in Eq. 2.
Alternative EGNNs such as NequIP [19], M3GNet [21], or MACE [20] which have been widely used for the development of MLIPs can also serve as plug-and-play encoders within OMG’s SI framework. Integrating different architectures is a direction that we plan to explore in future iterations of the framework.
IV.3 Comparison to other frameworks
We report results for the CSP task in Tab. 1 and for the DNG task in Tabs. 4 and 6 for DiffCSP, FlowMM, and OMG. We detail in Sec. V how we improve the extant benchmarks used in the field and therefore recompute all CSP and DNG benchmarks. In nearly all cases, we were able to generate structures using the DiffCSP and FlowMM source code which closely matched (within $\sim$1%) the previously reported metrics in their respective manuscripts. The observed differences can be attributed to the use of a newer version of SMACT composition rules [^2] The SMACT Python library updated its default oxidation states with the release of version 3.0. [43] and to natural fluctuations arising from model retraining.
Since the focus of this work is to assess our model’s ability to learn unconstrained and unconditioned flows, we do not compare against symmetry-constrained generation methods [48, 49, 50, 51, 52]. Symmetry constraints can be incorporated in future extensions of the flexible OMG framework.
V Experiments
@@
Table 2: Results from de novo generation of 10,000 structures with models trained on the MP-20 dataset. The integration steps for OMG is chosen based on best overall performance. For DiffCSP and FlowMM, we recomputed all metrics. All OMG models in this table use the ODE sampling scheme. Best scores in each category are bolded. [^4] We do not bold any values in the structural validity category as the CDVAE model reports the state-of-the-art with 100% structural validity. For the Wasserstein density and $N$ary metrics, we only bold values lower than 0.075 and 0.079 respectively, as these were the values reported by FlowMM for their model with 500 integration steps (not included in this table).
@@
OMG
Linear; $\gamma=0$ 150 99.64 87.02 86.77 99.39 99.86 0.0834 0.0784 0.1371
Trig; $\gamma=\sqrt{1.19t(1-t)}$ 690 97.89 85.74 83.97 98.51 99.83 0.0646 0.1914 0.1806
Enc-Dec; $\gamma=\sin^{2}{\pi t}$ 840 97.25 86.35 84.19 99.62 99.61 0.1155 0.0553 0.0465
SBDM; $\gamma=0$ 690 96.38 84.19 81.14 98.94 99.80 0.0788 0.1287 0.1127
CFP + CSP (Linear, $\gamma=0$) 130+210 97.95 79.68 78.21 99.67 99.50 0.5614 0.2008 0.6256
Table 3: Stability, uniqueness, and novelty results from de novo generation on the MP-20 dataset computed for the same models as in Tab. 4. All evaluations are performed with the MatterGen code base [53] with respect to the included reference Alex-MP-20 dataset. The average RMSD is between the generated and the relaxed structures, and the average energy above hull is reported in units of eV/atom.  [^6] OMG’s performance is comparable to the 22.56% S.U.N. Rate and 0.11Å RMSD reported by the MatterGen-MP model when trained on MP-20 [31], though we cannot directly compare to these benchmarks as those structures are relaxed using DFT.
Method $\langle E\rangle/N$ ($\downarrow$) RMSD (Å, $\downarrow$) Stability (%, $\uparrow$) S.U.N. (%, $\uparrow$)
above hull Rate Rate
DiffCSP 0.1984 1.295 43.04 19.00
FlowMM 0.2509 0.651 37.47 13.86
OMG
Linear 0.1438 0.294 69.11 18.58
Trig 0.1842 0.763 49.49 19.12
Enc-Dec 0.1699 0.390 58.56 17.59
SBDM 0.2215 0.759 41.04 20.16
CFP + CSP 0.2340 0.488 42.21 20.50
V.1 Benchmarks and datasets
We use the following datasets to benchmark the OMG model: perov-5 [54], a dataset of perovskites with 18,928 samples with five atoms per unit cell in which only lattice lengths and atomic types change; MP-20 [55, 29] from the Materials Project that contains 45,231 structures with a maximum of $N=20$ atoms per unit cell, and MPTS-52 [56] which is a time split of the Materials Project with 40,476 structures with up to $N=52$ atoms per unit cell and is typically the most difficult to learn. We use the same 60-20-20 splits as Xie et al. [29], Jiao et al. [30], Miller et al. [32]. For the Alex-MP-20 dataset [31], we used an 80-10-10 split constructed from MatterGen’s 90-10 split, in which we removed 10% of the training data to create a test dataset. This dataset contains 675,204 structures with 20 or fewer atoms per unit cell from the Alexandria [57, 58] and MP-20 datasets. We do not include the carbon-24 dataset [59] in our results, as match rate is ill-defined for this dataset; because all elements are carbon, it is not clear how many generated structures are unique and producing a structure that matches one in the reference dataset is trivial [^7] Previous papers [29, 30, 32] report match rate for carbon-24, but they do not compare the generated structure to the entirety of the reference dataset; their results suggest the match tolerance is larger than the differences between the carbon-24 structures..
V.2 Performance metrics
@@
We assess the state-of-the-art performance of OMG’s models using a variety of standard, refined, and contributed benchmarks. For CSP (see Tab. 1), the match rate and root mean square distance (RMSD) of matched structures are computed between the test dataset and the generated structures using Pymatgen’s StructureMatcher module [60] with tolerances ($\texttt{stol}=0.5$, $\texttt{ltol}=0.3$, $\texttt{angletol}=10$).
For DNG (see Tab. 4), metrics include validity (structural and compositional), coverage (recall and precision), and Wasserstein distances between distributions of properties including density $\rho$, number of unique elements $N$ (i.e., an $N$-ary material), and average coordination number by structure $\langle CN\rangle$. We introduce the average coordination number benchmark in part due to the difficulty of generating symmetric structures; a structure’s average coordination number is a useful fingerprint, and higher-coordinated structures tend to be more symmetric.
The previous DNG metrics are used during optimization of hyperparameters (see Appendix B.3), with the best models on the validation set being used to calculate stability rate as well as the S.U.N. (stable, unique, and novel) rate (see Tab. 6). The stability-based metrics are calculated using MatterGen’s code base [53], in which the machine-learned interatomic potential MatterSim [61] is utilized for structural relaxation. This requires significantly less compute time in comparison to running DFT relaxations. For all models with reported S.U.N. rates, we generated a set of 10,000 structures which are filtered to remove those with elements not handled by MatterSim. Further details are provided in Appendix C. In addition to the stability and S.U.N rates, we also provide in Tab. 4 the average RMSD between the generated and relaxed structures, as well as the average energy above the convex hull. Figure 3 compares the histograms of the energies above hull of the generated structures between OMG, DiffCSP, and FlowMM, and shows OMG’s superior performance for the generation of stable structures.
Figure 3: Histogram of the computed energies above the convex hull for structures generated by FlowMM, DiffCSP, and OMG (linear interpolant). The OMG model consistently produces lower energy structures compared to FlowMM and DiffCSP. See Appendix C.3 for calculation details.
@@
We have refined the calculation of several metrics, and reported these for all models. Previously reported match rates filtered the matched generated structures by their validity according to structural constraints on interatomic distances $r>0.5$Å as well as SMACT composition rules [43] which screen structures by their electron valence and charge neutrality. We note, however, that the datasets themselves contain invalid structures—for example, the MP-20 test dataset has $\sim 10$% compositionally invalid structures. Thus, we argue that the removal of these invalid structures for computation of match rate and RMSE is not reasonable for assessing learning performance; we do, however, provide both match rates (with and without validation filtering).

### v2

@@ III.3 Joint Generation with Stochastic Interpolants
For both CSP and DNG tasks, we seek to generate samples from a joint distribution over multiple coordinates. For DNG, this joint distribution $\rho_{1}$ encompasses all elements of a crystal unit cell. For CSP we similarly model the joint distribution, $\rho_{1}$, but with atom types fixed to compositions sampled from the target dataset. For both tasks, the total loss function is formulated as a weighted sum of the individual loss functions for each variable (see Appendix C.2), and their relative weights are optimized (see Appendix C.3). We illustrate both types of models and their structure generation process in Fig. 2a.
Figure 2: Illustration of crystal structure prediction (CSP) and de novo generation (DNG) tasks. (a) For CSP, the species $\bm{A}$ are fixed with known compositions from $t=0$. From this, we predict $\bm{X}$ and $\bm{L}$ from randomly sampled initial values. For DNG, we predict $(\bm{A},\bm{X},\bm{L})$ jointly. Our implementation of discrete flow matching (DFM) initializes $\bm{A}$ as a sequence of masked particles that are unmasked through a series of discrete jumps to reveal a physically reasonable composition. (b) Two avenues for performing DNG of materials. The first uses two steps: a chemical formula prediction (CFP) model predicts compositions and then uses a CSP model to find accompanying stable structures. The second trains a DNG model over cell, species, and fractional coordinates jointly as shown in (a).
Additionally, for DNG, we consider a two-step process in which composition is learned separately from structure, as seen in Fig. 2b. In this approach, we first train a chemical formula prediction (CFP) model (see Appendix C.1) to generate compositions optimized for SMACT stability [45], similarity in the distribution of $N$-arity of known structures, as well as uniqueness and novelty. The predicted compositions are then used as input for a pretrained CSP model, which generates the corresponding atomic configurations.
@@ IV.1 Choice of Interpolant
In training OMatG, we optimize the choice of the interpolating function that is used during training for the lattice vectors (without periodic boundary conditions) and the fractional coordinates (with periodic boundary conditions). We consider four interpolants of the form defined in Eq. (1), each shaping the interpolation trajectory differently (see also Appendices B.2 and B.9 for further details).
The linear interpolant defines a constant velocity trajectory from $x_{0}$ to $x_{1}$. When combined with an ODE sampling scheme and $\gamma=0$, this reproduces the specific instantiation of CFM implemented in FlowMM (see Appendix B.9). However, combining the linear interpolant with an SDE sampling scheme or nonzero $\gamma$ already introduces key differences. The inclusion of the latent variable can promote smoother learned flows [33], while stochastic sampling alters the generative dynamics compared to the deterministic formulation in FlowMM. The trigonometric interpolant prescribes trajectories with more curvature than the linear interpolant. The encoder-decoder interpolant first evolves samples from $\rho_{0}$ at $t=0$ to follow an intermediate Gaussian distribution at a switch time $T_{\text{switch}}$, before mapping them to samples from the target distribution $\rho_{1}$ at $t=1$. This approach has been found to interpolate more smoothly between distributions, potentially mitigating the formation of spurious features in the probability path at intermediate times [33]. Lastly, we consider variance-preserving score-based diffusion (VP SBD) and variance-exploding score-based diffusion (VE SBD) interpolants. When paired with an SDE sampling scheme, these interpolants are mathematically equivalent to the corresponding SBDM, but on the continuous time interval $[0,1]$. Different noise schedules of the variance-preserving and variance-exploding SBDMs can likewise be encoded in different variants of the VP and VE SBD interpolants (see Appendix B.9). For the results presented in this paper, we only consider a constant noise schedule for the VP SBD interpolant, and a geometric noise schedule for the VE SBD interpolant. The SBD interpolants assume that $\rho_{0}$ is a Gaussian distribution, and unlike the previous interpolants it involves no explicit latent variable; instead the $\alpha(t)x_{0}$ term takes on this role. The incorporation of VP and VE SBD interpolants enables OMatG to reproduce similar conditions to those in DiffCSP and MatterGen.
The trajectory of the encoder-decoder interpolant between times $t=T_{\text{switch}}$ and $t=1$ resembles that of the SBD interpolants between times $t=0$ and $t=1$. For the example of using the encoder-decoder interpolant only for the coordinates, however, we emphasize that the Gaussian-distributed coordinates at $t=T_{\text{switch}}$ are conditioned on other coordinates that are partially interpolated at this point. Conversely, for SBD interpolation, the Gaussian distributed coordinates at $t=0$ are only conditioned on other random variables since, at this point, all elements of $x_{0}$ are randomly distributed.
To investigate how different interpolants affect generative performance, we consider all interpolants outlined above for both the atomic positions $\bm{X}$ and the lattice vectors $\bm{L}$. We noted that learning accurate velocities and denoisers for the atomic positions, $\bm{X}$ is more challenging than for the other degrees of freedom. Accordingly, we optimize all hyperparameters—including the choice of interpolant for $\bm{L}$—separately for each interpolant applied to $\bm{X}$. This results in a set of experiments specific to the positional interpolants, where the best performing lattice interpolant may vary.
IV.2 Equivariant Representation of Crystal Structures
In the current OMatG implementation, we employ CSPNet as an encoder that is trained from scratch. The CSPNet architecture encodes atomic types using learnable atomic embeddings and represents fractional coordinates through sinusoidal positional encodings (see Appendix C.1). These features are processed through six layers of message-passing, after which the encoder produces the velocity $b^{\theta}(t,x)$ of both the lattice and the fractional coordinates, as well as potentially predicting the denoiser $z^{\theta}(t,x)$. For DNG, the network must also predict $\log p_{1|t}^{\theta}(\bm{a}_{1}|x_{t})$. The resulting outputs inherently preserve the permutation, rotational, and translational symmetries embedded in CSPNet.
The output of CSPNet is invariant with respect to translations of the fractional coordinates in the input. Thus, one should, in principle, use a representation of the fractional coordinates that does not contain any information about translations. While this is straightforward in Euclidean space by removing the mean of the coordinates of the given structure, this cannot be done with periodic boundary conditions where the mean is not uniquely defined. We follow Miller et al. [32] and instead remove the well-defined center-of-mass motion when computing the ground-truth velocity $\partial_{t}x(t,x_{0},x_{1},z)$ in Eq. (2).
Alternative EGNNs such as NequIP [19], M3GNet [21], or MACE [20] which have been widely used for the development of MLIPs can also serve as plug-and-play encoders within OMatG’s SI framework. Integrating different architectures is a direction that we plan to explore in future iterations of the framework.
IV.3 Comparison to other Frameworks
We compare our results to DiffCSP and FlowMM models for both CSP and DNG. For DNG, we additionally consider the MatterGen-MP model that was trained on the same MP-20 dataset as OMatG’s DNG models. We detail in Section V how we improve the extant benchmarks used in the field and therefore recompute all CSP and DNG benchmarks for these models. In nearly all cases, we were able to generate structures using the DiffCSP, FlowMM, and MatterGen source codes whose metrics closely matched the previously reported metrics in their respective manuscripts. For DiffCSP and MatterGen, we relied on published checkpoints while we retrained FlowMM from scratch. The observed differences can be attributed to the use of a newer version of SMACT composition rules [^3] The SMACT Python library updated its default oxidation states with the release of version 3.0. [45] and to natural fluctuations during generation and model retraining.
Since the focus of this work is to assess our model’s ability to learn unconstrained and unconditioned flows, we do not compare against symmetry-constrained generation methods [47, 48, 49, 50, 51]. Symmetry constraints can be incorporated in future extensions of the flexible OMatG framework.
V Experiments
V.1 Performance Metrics
@@
Table 1: Results from crystal structure prediction. Match rate and RMSE of matched structures without (left) and with (right) filtering for structural and compositional validity are reported for all models. For OMatG’s models, the choice of positional interpolant, latent variable component $\gamma$, and sampling scheme are noted. For all SDE sampling schemes the inclusion of $\gamma$ is assumed and not noted; for SBD interpolants $\gamma$ is not relevant. Further details and complete results for perov-5 and MP-20 can be found in Appendix C.3.
@@
OMatG
Linear (ODE) w/o $\gamma$ 51.86 / 50.62 0.0757 / 0.0760 69.83 / 63.75 0.0741 / 0.0720 27.38 / 25.15 0.1970 / 0.1931 72.02 / 64.23 0.0683 / 0.0671
Linear (SDE) w/ $\gamma$ 74.16 / 72.87 0.3307 / 0.3315 68.20 / 61.88 0.1632 / 0.1611 23.95 / 21.70 0.2402 / 0.2353 61.07 / 54.45 0.1870 / 0.1860
Trig (SDE) w/ $\gamma$ 73.37 / 71.60 0.3610 / 0.3614 68.90 / 62.65 0.1249 / 0.1235 24.51 / 22.26 0.1867 / 0.1804 72.50 / 64.71 0.1261 / 0.1251
Enc-Dec (ODE) w/ $\gamma$ 68.08 / 64.60 0.4005 / 0.4003 55.15 / 49.45 0.1306 / 0.1260 14.65 / 13.53 0.2543 / 0.2500 68.11 / 60.58 0.0957 / 0.0938
VP SBD (ODE) 83.06 / 81.27 0.3753 / 0.3755 45.57 / 39.48 0.1880 / 0.1775 9.66 / 8.36 0.3088 / 0.3041 46.23 / 39.96 0.1718 / 0.1618
VE SBD (ODE) 60.18 / 52.97 0.2510 / 0.2337 63.79 / 57.82 0.0809 / 0.0780 21.42 / 19.57 0.1740 / 0.1702 67.79 / 60.25 0.0674 / 0.0649
@@
Table 2: Results from de novo generation of 10 0001000010\,00010 000 structures with models trained on the MP-20 dataset. The integration steps for OMatG is chosen based on best overall performance. For OMatG’s model, the choice of positional interpolant, latent variable component $\gamma$, and sampling scheme are noted. Best scores in each category are bolded.
@@
MatterGen-MP 1000 99.93 83.89 83.89 96.62 99.90 0.2741 0.1632 0.4155
OMatG
Linear (SDE) w/ $\gamma$ 710 99.04 83.40 83.40 99.47 98.81 0.2583 0.0418 0.4066
Trig (ODE) w/ $\gamma$ 680 95.05 82.84 82.84 99.33 94.75 0.0607 0.0172 0.1650
Enc-Dec (ODE) w/ $\gamma$ 840 97.25 86.35 84.19 99.62 99.61 0.1155 0.0553 0.0465
VP SBD (SDE) 870 93.38 80.66 80.66 98.95 92.76 0.1865 0.0768 0.1637
CFP + CSP [Linear (ODE) w/o $\gamma$] 130+210 97.95 79.68 78.21 99.67 99.50 0.5614 0.2008 0.6256
We do not bold any values in the structural validity category as the CDVAE model reports the state of the art with 100% structural validity. For the Wasserstein distances of the density and $N$ary distributions, we only bold values lower than 0.075 and 0.079 respectively, as these were the values reported by FlowMM for their model with 500 integration steps (not included in this table).
@@
Crystal structure prediction. We assess the performance of OMatG’s and competing models using a variety of standard (introduced in [29, 31]), refined, and contributed benchmarks. For the CSP task, we generate a structure for every composition in the test dataset. We then attempt to match every generated structure with the corresponding test structure using Pymatgen’s StructureMatcher module [52] with tolerances ($\texttt{stol}=0.5$, $\texttt{ltol}=0.3$, $\texttt{angletol}=10$). We finally report the match rate and the average root-mean square displacement (RMSE) between the test structures and matched generated structures. Here, the RMSEs computed by Pymatgen are normalized by $(V/N)^{1/3}$, where $V$ is the (matched) volume and $N$ is the number of atoms. During hyperparameter optimization, we only attempt to maximize the match rate (see Appendix C.3).
@@
Previously reported match rates filtered the matched generated structures by their structural and compositional validity (see Appendix D.2). We note, however, that the datasets themselves contain invalid structures—for example, the MP-20 test dataset has $\sim 10$% compositionally invalid structures. Thus, we argue that the removal of these invalid structures for computation of match rate and RMSE is not reasonable for assessing learning performance; we do, however, provide both match rates (with and without validation filtering).
De novo generation. For the DNG task, metrics include validity (structural and compositional), coverage (recall and precision), and Wasserstein distances between distributions of properties including density $\rho$, number of unique elements $N$ (i.e., an $N$ary material), and average coordination number by structure $\langle CN\rangle$. We newly introduce the average coordination number benchmark due to the difficulty of generating symmetric structures; a structure’s average coordination number is a useful fingerprint, and higher-coordinated structures tend to be more symmetric.
The previous DNG metrics are used in conjunction during optimization of hyperparameters (see Appendix C.3). For the best models, we then structurally relax the generated structures in order to calculate the stability and the S.U.N. (stable, unique, and novel) rates using MatterGen’s code base [53]. The S.U.N. rate is defined as the percentage of generated structures that are stable with respect to a reference convex hull (within 0.1 eV/atom), are not found within the reference set (novel), and are not duplicated within the generated set itself (unique). The machine-learned interatomic potential MatterSim [54] is utilized for initial structural relaxation, which is subsequently followed by a more computationally expensive DFT relaxation. Full workflow details are given in Appendix D.4. In addition to the stability-based metrics, we also report the root mean squared displacement (RMSD) between the generated and relaxed structures (unnormalized, in units of Å).
V.2 Benchmarks and Datasets
We use the following datasets to benchmark the models: perov-5 [55], a dataset of perovskites with 18 9281892818\,92818 928 samples with five atoms per unit cell in which only lattice lengths and atomic types change; MP-20 [56, 29] from the Materials Project that contains 45 2314523145\,23145 231 structures with a maximum of $N=20$ atoms per unit cell, and MPTS-52 [57] which is a chronological data split of the Materials Project with 40 4764047640\,47640 476 structures with up to $N=52$ atoms per unit cell and is typically the most difficult to learn. We use the same 60-20-20 splits as Xie et al. [29], Jiao et al. [30], Miller et al. [32]. Additionally, we consider the Alex-MP-20 dataset [31], where we used an 80-10-10 split constructed from MatterGen’s 90-10 split, in which we removed 10% of the training data to create a test dataset. This dataset contains 675 204675204675\,204675 204 structures with 20 or fewer atoms per unit cell from the Alexandria [58, 59] and MP-20 datasets. We do not include the carbon-24 dataset [60] in our results, as the current match rate metric is ill-defined for this dataset; because all elements are carbon, it is not clear how many generated structures are unique and producing a structure that matches one in the reference dataset is trivial. [^4] Previous papers [29, 30, 32] report match rate for carbon-24, but they do not compare each generated structure to the entirety of the reference dataset; their results suggest the match tolerance is larger than the differences between the carbon-24 structures themselves.
V.3 Results
Crystal structure prediction. We report the CSP performance of DiffCSP, FlowMM, and six OMatG models on the four benchmark datasets in Tab. 1. Further ablation results across OMatG variants can be found in Tables 6 and 8 in the Appendix. OMatG significantly outperforms previous approaches with respect to match rate on all datasets. We highlight the strong match rates on the perov-5 dataset achieved using the VP SBD and trigonometric positional interpolants with ODE sampling schemes, as shown in Table 1 and Table 6, that greatly surpass (by a factor of $1.6$) the match rates of previous models. We note that the relative performance for ODE vs. SDE sampling schemes depends on the positional interpolant. OMatG also outperforms previous models’ match rate for CSP on the MP-20 and MPTS-52 datasets with the linear (both ODE and SDE sampling scheme) and the trigonometric positional interpolants. Finally, OMatG establishes the first performance baseline for the CSP task on the Alex-MP-20 dataset.
De novo generation. For DNG models, we show the validity, coverage, and property metrics of the models in Tab. 2 and the stability, uniqueness, and novelty results in Tab. 3. Further ablation results across OMatG variants can be found in Tab. 10 in the Appendix, and qualitative plots of various distributions are compared in Appendix D.5. OMatG achieves state-of-the-art performance over DiffCSP, FlowMM, and MatterGen-MP for multiple positional interpolants thanks to the broader design space brought by the SIs. Figure 3 compares the distributions of the average energies above the hull for generated structures, exhibiting OMatG’s superior performance for the generation of stable structures. OMatG consistently produces lower energy structures compared to previous models, and they are also generated close to their relaxed configuration. This, together with high novelty rates, begets improved S.U.N. rates. OMatG also outperforms FlowMM in settings where large language models are used as base distributions [61], with results shown in Appendix E.
Table 3: Stability (defined as $\leq 0.1$ eV/atom above hull), uniqueness, and novelty results from de novo generation on the MP-20 dataset computed for the same models as in Tab. 2. All evaluations are calculated with respect to $1000$ relaxed structures (see Appendix D.4), utilizing the MatterGen code base [53] and the included reference Alex-MP-20 dataset. The average RMSD is between the generated and the relaxed structures, and the average energy above hull is reported in units of eV/atom.
Method $\langle E\rangle/N$ ($\downarrow$) RMSD Novelty Stability S.U.N.
above hull (Å, $\downarrow$) Rate (%, $\uparrow$) Rate (%, $\uparrow$) Rate (%, $\uparrow$)
DiffCSP 0.1751 0.3861 70.04 43.43 15.95
FlowMM 0.1917 0.6316 69.13 41.20 11.73
MatterGen-MP 0.1772 0.1038 72.40 44.79 20.30
OMatG
Linear (SDE) w/ $\gamma$ 0.1808 0.6357 73.31 46.18 22.48
Trig (ODE) w/ $\gamma$ 0.1670 0.6877 66.45 52.81 21.10
Enc-Dec (ODE) w/ $\gamma$ 0.1482 0.4187 55.21 60.04 18.77
VP SBD (SDE) 0.2120 0.7851 76.25 40.83 20.73
CFP + CSP 0.2302 0.5375 79.08 39.38 20.17
Figure 3: Histogram of the computed energies above the convex hull for structures generated by FlowMM, DiffCSP, MatterGen-MP, and OMatG (Enc-Dec interpolant). The OMatG model consistently produces lower energy structures compared to competing models. See Appendix D.4 for calculation details.
V.4 Discussion
Crystal structure prediction. We generally observe during hyperparameter optimization that the main learning challenge lies in the accurate prediction of the atomic coordinates, which tended to have a higher relative weight in calculating the full loss function in Eq. (32) (see Tabs 7 and 9 in the Appendix). However, the two best-performing models for the perov-5 dataset instead exhibited the opposite—lending the most weight to learning of the cell vectors.
Our CSP results indicate a tradeoff between match rate and RMSE—which is only computed if matched. We find that as the number of generated structures matching known compositions in the test dataset increases, the structural fidelity—quantified by the accompanying RMSE—also tends to increase. This tradeoff most strongly influences our results on the perov-5 dataset where all positional interpolants but the linear one can beat the previous state-of-the-art match rate. Here, particles generally find the correct local chemical configurations to flow towards during generation, but are not able to end up in the precise symmetric sites. The ODE-based linear interpolant without a latent variable, in contrast, has the lowest RMSE because the particles flow to more symmetric positions, but the local environments are not correct due to species mismatch. We quantify this effect in Fig. 7 in the Appendix.
For the CSP task on the perov-5 dataset, we highlighted the particularly strong performance of the VP SBD and trigonometric interpolants in achieving a high match rate. Unlike other datasets, perov-5 has a fixed number of $N=5$ atoms per unit cell and a fixed (cubic) cell with varying side lengths and similar fractional positions—a combination which should not expose the model to a large variety of unit cell choices during interpolation or generation. By contrast, in other datasets, no singular representation of the periodic repeat unit is imposed on flows, meaning the model cannot learn the invariance (or even equivariance) to the choice of periodic repeat unit. [^5] Using Niggli reduction during learning to enforce a unique choice of unit cell on structures from our datasets is not sufficient for enforcing this invariance during generation of structures. This likely contributes to the difficulty of unconstrained flow-based models in generating highly symmetric structures. Thus, the perov-5 dataset presents a unique case where the invariance to unit cell choice does not need to be learned, making this dataset a useful benchmark for evaluating positional interpolant performance. It is possible that the superior performance of the VP SBD and trigonometric interpolants arise from their ability to generate more circuitous flow trajectories compared to the strictly geodesic paths imposed by the linear interpolant—akin to the reasoning behind using latent variables to enhance learning in SIs [33].
De novo generation. Thorough hyperparameter optimization enables us to note trends among the best-performing DNG models. In Tabs 10 and 11 in the Appendix, we show the performance metrics and hyperparameters for each model by choice of positional interpolant, sampling scheme, and $\gamma(t)$ in the latent variable. We observe that several of our best performing models (with respect to S.U.N. and RMSD) possess lower levels of ‘species noise’ $\eta$ which sets the probability that an atom will change its identity if already in an unmasked state (see Appendix B.6). Additionally, we find that linear and trigonometric interpolants favor an element-order permutation as a data-dependent coupling during training (see Appendix B.8), while the encoder-decoder and SBD interpolants prefer to not use this coupling. Finally, we note that VP SBD models require a similar magnitude of velocity annealing during generation for positions and lattices (see Appendix B.7). This is in stark contrast to all other models, where a significantly larger velocity annealing parameter is required for generating the positions.